OpenStudy (dallaslifebaby01):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (mortonsalt):

Here is how your sequence will go: n=-10 The next n would have a difference of -2 from the first n. n_2 = -10-2=-12 The third n would have a difference of -2 from the second n. n_3 = -12-2 = -14 And so on and so forth until you reach the tenth term.

OpenStudy (dallaslifebaby01):

okay thats what i was doing i was worried i was doing it wrong